A Pumpkin gets launched into the air from a cliff above a pond. The following equation represents the height of the pumpkin from the pond, h(t)=−16t2+56t+32where t is time in seconds.  What is the maximum height the pumpkin reaches?

1
Expert's answer
2022-05-13T11:43:04-0400

"h(t)=-16t^2+56t+32"

"h'(t)=-32t+56=0"

"t=56\/32=1.75"

"h''(t)=-32<0"

So it is maximum

"h(t)=-16(1.75^2)+56(1.75)+32=81"